OpenStudy (anonymous):

Find the requested probability. P(A) if P(A) = 0.7

OpenStudy (jmark):

Given P(A) = 0.7 then =>P(A) = 1 - P(A) = 1 - 0.7 = 0.3 it is also written as P(A') or P(Ac) Learn more about Probability http://goo.gl/zyv84d